OpenStudy (anonymous):

through how many radians does the second hand of a watch rotate in 3 minutes

OpenStudy (kropot72):

In one rotation the second hand goes though -2pi radians.

OpenStudy (anonymous):

so what is the equatoin?

OpenStudy (kropot72):

How many seconds are there in one minute?

OpenStudy (anonymous):

60

OpenStudy (anonymous):

oh lol i got it. its 6pi. ok how about .... throught how many degrees does the minute hand of a clock rotate in 3 minutes.

OpenStudy (kropot72):

Correct. So the required number of radians is \[-2\times \pi \times 60\times 3\]

OpenStudy (anonymous):

thank you

OpenStudy (kropot72):

You're welcome.

OpenStudy (mathstudent55):

How many rotations does the seconds hand finish in 1 minute? How many rotations does the seconds hand fisnish in 3 minutes? How many degrees are there in one full rotation? How many degrees are there in 3 full rotations?